3.) Data on 85 roller coasters is collected on the duration of the ride in seconds and the height of the initial drop in feet. The regression analysis gives the model Duration= 85.59+0.369 Drop.

A new roller coaster advertises an initial drop of 200 feet. How long would you predict the rides​ last?
Answer #1

The regression equation is given by ,

Model Duration = 85.59 + 0.369 Drop

Since , Drop = 200 feet

Therefore ,

Model Duration = 85.59 + 0.369 Drop = 85.59 + 0.369*200 = 159.39 Seconds